# D-Judge: How Far Are We? Accessing the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance

### v2

@@ Abstract
In Artificial Intelligence Generated Content (AIGC), distinguishing AI-synthesized images from natural ones remains a key challenge. Despite advancements in generative models, significant discrepancies persist. To systematically investigate and quantify these discrepancies, we introduce an AI-Natural Image Discrepancy accessing benchmark (D-Judge) aimed at addressing the critical question: how far are AI-generated images (AIGIs) from truly realistic images? We construct D-ANI, a dataset with 5,000 natural images and over 440,000 AIGIs generated by nine models using Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I) prompts. Our framework evaluates the discrepancy across five dimensions: naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and human validation. Results reveal notable gaps, emphasizing the importance of aligning metrics with human judgment. Source code and datasets are available at https://shorturl.at/l83W2.
Artificial Intelligence Generated Content, AI-synthesized Images, Multimodal, Discrepancy, Evaluation Framework
1 Introduction
@@
With the rapid advancement of deep learning techniques, the proliferation of Artificial Intelligence Generated Content (AIGC) has garnered significant attention across various domains, such as e-commerce, gaming, medicine, animation, and autonomous driving (Li et al., 2024; Qian et al., 2024). AI-generated Images (AIGI) have been one of the mainstream forms of AIGC, and a variety of AI image generative models have been proposed to make the generated synthetic images as realistic as natural images, ranging from the earlier generative adversarial networks (GANs) (Tao et al., 2023, 2022), advanced diffusion models (DMs) (Xu et al., 2023b; Wei et al., 2023) to the large multimodal generative model like DALL·E (Ramesh et al., 2022).
Figure 1: Toy example demonstrating the key aspects of D-Judge.
@@
Despite the rapid development of AI image generative models, AI-generated images still fall short of real-world application standards due to discrepancies with realistic natural images (Li et al., 2023a). To address this, various AI-generated image quality assessment methods have emerged (Wang et al., 2023; Xu et al., 2023a; Hu et al., 2023). For instance, Wang et al. (Wang et al., 2023) introduced the AIGCIQA2023 database and a subjective evaluation framework based on quality, authenticity, and text-image correspondence. Similarly, PKU-I2IQA (Yuan et al., 2023) developed a perception-based database for image-to-image generation using no-reference and full-reference methods. Other studies, such as Pick-a-Pic (Kirstain et al., 2023), HPS v2 (Wu et al., 2023), TIFA (Hu et al., 2023), and ImageReward (Xu et al., 2023a), focused on training unified score models for automatic AIGC image quality assessment, moving beyond traditional metrics like Inception Score (Salimans et al., 2016) and FID (Heusel et al., 2017).
@@
To address the above challenges, we proposed an AI-Natural Image Discrepancy Evaluation Benchmark (D-Judge) to investigate and interpret what discrepancies still remained between AI-generated images and natural images, finally answering the important question “How far are AI generative models with respect to the visual forms of images?”. Especially, we have two important contributions with reference to the key problems. Large Multimodal Evaluation Dataset construction: we have curated a large multimodal evaluation dataset called D-ANI (Distinguishing Natural and AI-generated) dataset, with about 440,000 AIGI generated from 9 representative generative models based on both unimodal and mutimodal guidance including Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I). The data size of our DNAI dataset scales to 100x that of prior datasets. Fine-grained Evaluation Framework, we propose a fine-grained evaluation framework to conduct the systematic and comprehensive assessment and evaluation of DNAI, covering 5 diverse and important aspects, including naive visual feature quality, semantic alignment among multimodal generation, aesthetic appeal, downstream applicability, and the coordinated human validation. The toy example is illustrated in Figure 1. For example, given a text prompt, an image is generated using a generative model and paired with a natural reference image retrieved from the natural database. D-Judge then evaluates the discrepancy rate between the two images across the aforementioned dimensions.
@@ 2 Related Work
Over the past decades, numerous image quality assessment methods have been developed to evaluate natural images (N. et al., 2015). These methods have focused on various visual features and properties, such as perceptual appearance (Zhang et al., 2018), naturalness (Ma et al., 2018), and aesthetics (Esfandarani & Milanfar, 2018). For instance, BRISQUE evaluates natural image quality by analyzing spatial domain features, while PIQE assesses perceptual quality through block-based image segmentation, both serving as effective no-reference metrics. Other notable measures include FID and Inception Scores, which assess the quality and diversity of generated images, and SSIM (Wang et al., 2004) and PSNR, which quantify structural similarity and pixel-level accuracy. Beyond quantitative measures, human evaluations have also been leveraged to estimate natural image quality. For example, NIMA (Esfandarani & Milanfar, 2018), proposed by Talebi and Milanfar, uses a deep CNN trained on human-rated images to predict aesthetic quality. Similarly, Wong et al. introduced the AVA dataset, which facilitates aesthetic assessment based on human preferences (Murray et al., 2012).
@@
Recently, AI-generated content (AIGC) has seen significant advancements with the rise of generative models. Several methods and benchmarks have been introduced to evaluate AI-generated images (Zhang et al., 2023; Hu et al., 2023). These methods primarily focus on three aspects: perceptual quality, text-image correspondence, and aesthetics (Xu et al., 2023a; Wang et al., 2023), often training unified models to assess the overall quality of AI-generated images. For example, ImageReward (Xu et al., 2023a) presents a framework for aligning image generation with human preferences, AGIQA-3k (Li et al., 2023a) offers a comprehensive benchmark for assessing AI-generated image quality, and QBench (Wu et al., 2024a) evaluates content quality across multiple dimensions.
@@
3 D-Judge
Figure 2: The D-ANI dataset of D-Judge, guided by three types of prompts. Each generated image is paired with both the corresponding text and image from the naturally collected dataset (e.g., COCO).
Figure 3: The framework of D-Judge evaluates the Discrepancy Rate ($DR$) between AI-generated images and natural reference images across five dimensions: ➊ Naive Quality, ➋ Alignment, ➌ Aesthetic, and ➍ Applicability, complemented by ➎ human validation. The images with green borders represent reference natural images, while those with yellow borders are AI-generated counterparts. This comprehensive evaluation integrates pixel-level, frame-level, and distribution-level quality checks, alongside advanced tools like OpenAI CLIP, OpenAI Laion Aesthetic Evaluator, Qalign, ViT, U2Net, and BLIP2 to robustly assess the gap between the two image types. Where $CS$ means CLIP Score for the text and given images, and $AES$ means Aesthetic score.
@@ 3 D-Judge
3.1 Distinguishing AI-Natural Image Dataset
We construct the Distinguished Natural and AI-generated Image Dataset based on the classical natural image dataset MS COCO dataset. We select 5,000 images covering diverse captions from the MS COCO dataset and collect a total of 25,000 text-image pairs by pairing each image with five different pieces of text descriptions. The 25,000 text-image pairs are seen as the referenced natural image set.
For each text-image sample in the referenced natural image set, we collect AI-generated images from 9 representative generative models, guided by three types of prompts, including text-only prompts, image-only prompts, and text-image prompts. The 9 representative generative models are from diverse generative models ranging from the earlier generative adversarial networks (GANs) (e.g. GALIP (Tao et al., 2023), DF-GAN (Tao et al., 2022)) to the recent diffusion models (DMs) like Stable Diffusion v1.4 (SD14), v1.5 (SD15), v2.1 (SD21), XL (SDxl) (Rombach et al., 2022), Versatile-Diffusion (VD) (Xu et al., 2023b), OpenAI DALL·E 2 (D·E 2) (Ramesh et al., 2022) and DALL·E 3 (D·E 3) (OpenAI, 2024). The referenced natural images and the generated images comprise our distinguishing AI-Natural image dataset. More details about data construction and collection are presented in Appendix Sec. B.
Table 1 shows the statistics and properties of our D-ANI Dataset compared with the existing evaluation dataset. Our D-ANI Dataset has a total of 440,000 AI-generated images from 9 representative generated models. For details, please refer to Appendix Sec. B. Compared with state-of-the-art evaluation datasets, our D-ANI Dataset has convincing contributions:
  • D-ANI Dataset is the most extensive dataset for AI-generated image evaluation, and the data size scales to 100x.
  • D-ANI Dataset is guided by both unimodal prompts and multi-modal prompts (text-image). Prior datasets barely considered text prompts.
  • Comprehensive and Diverse generative models are exploited, ranging from the earlier GAN models, the recent popular diffusion models, and the influential large commercial generative model DALL·E 2 and DALL·E 3 issued by OpenAI.
Our large and comprehensive dataset works as the foundation of the assessment benchmark and enables thorough evaluation and analysis of the differences between AI-generated images and realistic natural images.
Table 1: Comparison of D-ANI dataset and Counterparts.
@@ 3.1 Distinguishing AI-Natural Image Dataset
AGIQA-1K 2 1,080 - - 1,080
AGIQA-3K 6 300 - - 2,982
AIGCIQA2023 6 100 - - 2,400
PKU-I2IQA 2 200 200 - 1,600
D-ANI (Ours) 9 25,000 5,000 25,000 445,000
3.2 D-Judge Framework
@@ 1. Naive Image Quality Discrepancy:
We leverage traditional image quality assessment methods for natural images to measure the AI-Natural image differences. In order to perform fine-grained analysis and exploration, we estimate the AI-Natural image differences from three levels by considering the inherent properties of image-style content, ranging from the pixel-level visual features, frame-level visual features to the holistic content distribution.
  • First, we use the full-referenced pixel-level similarity measures like SSIM, LPIPS, and DISTS to compute the visual similarities between the paired AI-generated image and natural image and use these calculated values to quantify the AI-Natural image differences from the low-level visual view.
  • Second, we use non-referenced frame-level visual features to qualify the perceptional quality of images and compute the frame-level visual quality difference rates between AI-generated images and natural images from the high-level visual view. The exploited frame-level visual features include PIQE (N. et al., 2015), IL-NIQE (Zhang et al., 2015), MUSIQ (Ke et al., 2021), DBCNN (Zhang et al., 2020), LIQE (Ma et al., 2018), Inception Score (Salimans et al., 2016), CLIPIQA (Radford et al., 2021b), TReS (Gu et al., 2015), HyperIQA (Su et al., 2020), UNIQUE (Zhang et al., 2021), BRISQUE , NIQE (Mittal et al., 2013), NRQM (Ma et al., 2018), etc. The Discrepancy Rate ($DR$) of these metrics provides insights into frame-level visual quality differences.
  • Third, we investigate the full-reference visual content distribution to capture the holistic differences between AI-generated images and natural images by computing the FID, CLIP-FID, and Inception Score of the AI-generated image set and the referenced natural image set. The results of these set-level measures can reflect the difference between AIGIs and natural ones from an integral perspective.
Based on the different levels of image quality measures, our framework can assess and measure the fine-grained AI-Natural images differences with Full-reference and Non-reference angles, covering both low-level and high-level. More details about the utilized traditional image quality assessment measures are described in Appendix Sec. C.
@@ 2. Semantic Alignment Discrepancy:
Generative models require the referenced prompts as semantic guidance to generate images. Therefore, semantic alignment can be an important quality indicator of the AIGIs. We use the widely used CLIP model (Radford et al., 2021a) to calculate the difference rates of Clip Scores, $DR_{CS}$ to represent alignment discrepancies between AIGIs and natural images, and provide a fine-grained analysis with respect to both unimodal and multimodal guidance.

### v3

# D-Judge: How Far Are We? Accessing the Discrepancies Between AI-synthesized and Natural Images through Multimodal Guidance

@@ Abstract.
In the rapidly evolving field of Artificial Intelligence Generated Content (AIGC), a central challenge is distinguishing AI-synthesized images from natural ones. Despite the impressive capabilities of advanced generative models in producing visually compelling images, significant discrepancies remain when compared to natural images. To systematically investigate and quantify these differences, we construct a large-scale multimodal dataset, D-ANI, comprising 5,000 natural images and over 440,000 AIGI samples generated by nine representative models using both unimodal and multimodal prompts, including Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I). We then introduce an AI–Natural Image Discrepancy assessment benchmark (D-Judge) to address the critical question: how far are AI-generated images (AIGIs) from truly realistic images? Our fine-grained evaluation framework assesses the D-ANI dataset across five dimensions: naive visual quality, semantic alignment, aesthetic appeal, downstream task applicability, and coordinated human validation. Extensive experiments reveal substantial discrepancies across these dimensions, highlighting the importance of aligning quantitative metrics with human judgment to achieve a comprehensive understanding of AI-generated image quality. The code [^1] https://github.com/ryliu68/DJudge and data [^2] https://huggingface.co/datasets/Renyang/DANI are publicly available.
Discrepancy Evaluation Benchmark, Image Naturalness Assessment, Multimodal Guided Generation, AI-generated images, AIGC
copyright: acmlicensedjournalyear: 2025copyright: ccdoi: 10.1145/3746027.3754837conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer visionccs: Computing methodologies Neural networksccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
1. Introduction
@@
With the rapid advancement of deep learning techniques, the proliferation of Artificial Intelligence Generated Content (AIGC) has garnered significant attention across various domains, such as e-commerce, gaming, medicine, animation, and autonomous driving (Li et al., 2024; Qian et al., 2024). AI-generated Images (AIGI) have been one of the mainstream forms of AIGC, and a variety of AI image generative models have been proposed to make the generated synthetic images as realistic as natural images, ranging from the earlier generative adversarial networks (GANs) (Tao et al., 2023, 2022), advanced diffusion models (DMs) (Xu et al., 2023; Wei et al., 2023) to the large multimodal generative model like DALL·E (Ramesh et al., 2022).
@@
Despite the rapid development of AI image generative models, AI-generated images still fall short of real-world application standards due to discrepancies with realistic natural images (Li et al., 2024). To address this, various AI-generated image quality assessment methods have emerged (Yuan et al., 2023; Xu et al., 2023; Hu et al., 2023). For instance, Wang et al. (Yuan et al., 2023) introduced the AIGCIQA2023 database and a subjective evaluation framework based on quality, authenticity, and text-image correspondence. Similarly, PKU-I2IQA (Yuan et al., 2023) developed a perception-based database for image-to-image generation using no-reference and full-reference methods. Other studies, such as Pick-a-Pic (Kirstain et al., 2023), HPS v2 (Wu et al., 2023), TIFA (Hu et al., 2023), and ImageReward (Xu et al., 2023), focused on training unified score models for automatic AIGC image quality assessment, moving beyond traditional metrics like Inception Score (Salimans et al., 2016) and FID (Heusel et al., 2017).
@@
Figure 1. Toy example demonstrating the key aspects of D-Judge.
To address the above challenges, we proposed an AI-Natural Image Discrepancy Evaluation Benchmark (D-Judge) to investigate and interpret what discrepancies still remained between AI-generated images and natural images, finally answering the important question “How far are AI generative models with respect to the visual forms of images?”. Especially, we have two important contributions with reference to the key problems. Large Multimodal Evaluation Dataset construction: we have curated a large multimodal evaluation dataset called D-ANI (Distinguishing Natural and AI-generated) dataset, with about 440,000 AIGI generated from 9 representative generative models based on both unimodal and mutimodal guidance including Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I). The data size of our DNAI dataset scales to 100x that of prior datasets. Fine-grained Evaluation Framework, we propose a fine-grained evaluation framework to conduct the systematic and comprehensive assessment and evaluation of DNAI, covering 5 diverse and important aspects, including naive visual feature quality, semantic alignment among multimodal generation, aesthetic appeal, downstream applicability, and coordinated human validation. The toy example is illustrated in Figure 1. For example, given a text prompt, an image is generated using a generative model and paired with a natural reference image retrieved from the natural database. D-Judge then evaluates the discrepancy rate between the two images across the aforementioned dimensions.
@@ 2. Related Work
Over the past decades, numerous image quality assessment methods have been developed to evaluate natural images (N. et al., 2015). These methods have focused on various visual features and properties, such as perceptual appearance (Zhang et al., 2018), naturalness (Ma et al., 2018), and aesthetics (Esfandarani and Milanfar, 2018). For instance, BRISQUE evaluates natural image quality by analyzing spatial domain features, while PIQE assesses perceptual quality through block-based image segmentation, both serving as effective no-reference metrics. Other notable measures include FID and Inception Scores, which assess the quality and diversity of generated images, and SSIM (Wang et al., 2004) and PSNR, which quantify structural similarity and structure-level accuracy. Beyond quantitative measures, human evaluations have also been leveraged to estimate natural image quality. For example, NIMA (Esfandarani and Milanfar, 2018), proposed by Talebi and Milanfar, uses a deep CNN trained on human-rated images to predict aesthetic quality. Similarly, Wong et al. introduced the AVA dataset, which facilitates aesthetic assessment based on human preferences (Murray et al., 2012).
@@
Recently, AI-generated content (AIGC) has seen significant advancements with the rise of generative models. Several methods and benchmarks have been introduced to evaluate AI-generated images (Zhang et al., 2023; Hu et al., 2023). These methods primarily focus on three aspects: perceptual quality, text-image correspondence, and aesthetics (Xu et al., 2023; Yuan et al., 2023; Li et al., 2024, 2025; Chen et al., 2025), often training unified models to assess the overall quality of AI-generated images. For example, ImageReward (Xu et al., 2023) presents a framework for aligning image generation with human preferences, AGIQA-3k (Li et al., 2024) offers a comprehensive benchmark for assessing AI-generated image quality, and QBench (Wu et al., 2024) evaluates content quality across multiple dimensions.
@@
3. D-Judge
@@
3.1. D-ANI Dataset
Figure 2. The D-ANI dataset of D-Judge, guided by three types of prompts. Each generated image is paired with both the corresponding text and image from the naturally collected dataset (e.g., COCO).
3.1.1. Natural Images
We constructed the Distinguished Natural and AI-generated Image dataset (D-ANI) based on the classical MS COCO natural image dataset (Lin et al., 2014). Specifically, we selected 5,000 carefully curated images from the MS COCO validation set, each accompanied by at least five unique captions, resulting in a total of 25,000 distinct text-image pairs. This rich textual diversity enhances the dataset’s applicability across various image-to-text and text-to-image tasks, making it an invaluable resource for evaluating and training image generation and understanding models. The 25,000 text-image pairs form the referenced natural image set, guiding generative models in producing new images in our study.
3.1.2. Generated Images
We collected AI-generated images from nine representative generative models using three distinct types of prompts: text-only (Text-to-Image, T2I), image-only (Image-to-Image, I2I), and combined text-image (Text-and-Image-to-Image, TI2I). These models range from earlier generative adversarial networks (GANs) like GALIP (Tao et al., 2023) and DF-GAN (Tao et al., 2022), to recent diffusion models (DMs) such as Stable Diffusion (versions v1.4, v1.5, v2.1, and XL) (Rombach et al., 2022), Versatile Diffusion (VD) (Xu et al., 2023), and OpenAI’s commercial models DALL·E 2 (D·E 2) (Ramesh et al., 2022) and DALL·E 3 (D·E 3) (OpenAI, 2024), which generate images via OpenAI’s API, while other models generate images locally. Detailed information w.r.t these models can be found in Appendix B.
Each model generated 25,000 images per guidance mode, except for DALL·E 2 and DALL·E 3, which produced 5,000 images per mode, leading to a total of 440,000 images. Due to intrinsic guidance capabilities, DF-GAN, GALIP, and DALL·E 3 were exclusively used in T2I mode, whereas DALL·E 2 operated in both T2I and I2I modes. The details are shown in Table 1.
The image resolution varies across these models: GALIP produces images at 224x224 pixels, DF-GAN at 256x256 pixels, models such as various versions of Stable Diffusion and DALL·E 2 at 512x512 pixels, and DALL·E 3 provides the highest resolution of 1024x1024 pixels. These generated images, combined with the referenced natural images, constitute our distinguishing AI-Natural image dataset.
Table 1. The details of the D-ANI dataset.
Types DFGAN GALIP SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 25,000 25,000 25,000 25,000 25,000 25,000 25,000 5,000 5,000
I2I - - 25,000 25,000 25,000 25,000 25,000 5,000 -
TI2I - - 25,000 25,000 25,000 25,000 25,000 - -
Table 2. Comparison of D-ANI dataset and Counterparts.
@@ 3.1.2. Generated Images
AGIQA-1K (Zhang et al., 2023) 2 1,080 - - 1,080
AGIQA-3K (Li et al., 2024) 6 300 - - 2,982
AGIQA-20K (Li et al., 2024) 15 200,000 - - 200,000
AIGCIQA2023 (Wang et al., 2023) 6 100 - - 2,400
ImageReward (Xu et al., 2023) 3 8,878 - - 136,892
HPDv2 (Wu et al., 2023) 9 108,000 - - 381,304
PKU-I2IQA (Yuan et al., 2023) 2 200 200 - 1,600
D-ANI (Ours) 9 25,000 5,000 25,000 440,000
Table 2 presents the key statistics and characteristics of our proposed D-ANI dataset in comparison with existing AIGC evaluation datasets. Specifically, the D-ANI dataset consists of 440,000 AI-generated images produced by nine representative generative models, alongside 25,000 human-annotated image-text pairs from the MS COCO validation set. This results in a total of 445,000 high-quality text-image pairs, supporting both generation and evaluation tasks. In contrast to state-of-the-art benchmarks, the D-ANI dataset offers several notable advantages:
  • It represents the most comprehensive dataset to date for evaluating AI-generated images, expanding the data volume by approximately 100 times compared to existing benchmarks.
  • It uniquely integrates prompts from both unimodal (text-only or image-only) and multi-modal (text-image combined) inputs, effectively addressing the prevalent limitation of missing multi-modal context in prior datasets.
  • It encompasses a broad and diverse range of generative models, spanning from early GAN-based architectures to the latest state-of-the-art diffusion techniques, including high-impact commercial systems such as OpenAI’s DALL·E 2 and DALL·E 3.
The extensive scale and diversity of the D-ANI dataset provide a robust foundation for systematically analyzing the differences between natural and AI-generated images under various conditions and prompt modalities. By encompassing a wide spectrum of image styles, resolutions, semantic content, and textual descriptions, the dataset facilitates rigorous and holistic evaluations. This, in turn, supports the development and benchmarking of more precise, robust, and trustworthy AI image generation systems for real-world deployment.
Figure 3. The framework of D-Judge evaluates the Discrepancy Rate ($DR$) between AI-generated images and natural reference images across five dimensions: ➊ Naive Quality, ➋ Alignment, ➌ Aesthetic, and ➍ Applicability, complemented by ➎ human validation. The images with green borders represent reference natural images, while those with yellow borders are AI-generated counterparts. This comprehensive evaluation integrates structure-level, frame-level, and distribution-level quality checks, alongside advanced tools like OpenAI CLIP, OpenAI Laion Aesthetic Evaluator, QAlign, ViT, U2Net, and BLIP2 to robustly assess the gap between the two image types. Where $CS$ means CLIP Score for the text and given images, and $AES$ means Aesthetic score.
3.2. D-Judge Framework
@@ 1. Naive Image Quality Discrepancy:
We leverage traditional image quality assessment methods for natural images to measure the AI-Natural image differences. In order to perform fine-grained analysis and exploration, we estimate the AI-Natural image differences from three levels by considering the inherent properties of image-style content, ranging from the structure-level visual features, frame-level visual features to the holistic content distribution.
  • First, we use the full-referenced structure-level similarity measures like SSIM, LPIPS, and DISTS to compute the visual similarities between the paired AI-generated image and natural image and use these calculated values to quantify the AI-Natural image differences from the low-level visual view.
  • Second, we use non-referenced frame-level visual features to qualify the perceptional quality of images and compute the frame-level visual quality difference rates between AI-generated images and natural images from the high-level visual view. The exploited frame-level visual features include PIQE (N. et al., 2015), IL-NIQE (Zhang et al., 2015), MUSIQ (Ke et al., 2021), DBCNN (Zhang et al., 2020), CLIPIQA (Radford et al., 2021), CNNIQA (Kang et al., 2014), TReS (Gu et al., 2015), HyperIQA (Su et al., 2020), LIQE (Ma et al., 2018), UNIQUE (Zhang et al., 2021), BRISQUE (Mittal et al., 2012) , NIQE (Mittal et al., 2013), NRQM (Ma et al., 2018), QAlign (Wu et al., 2024), Inception Score (IS) (Salimans et al., 2016), etc. The Discrepancy Rate ($DR$) of these metrics provides insights into frame-level visual quality differences.
  • Third, we investigate the full-reference visual content distribution to capture the holistic differences between AI-generated images and natural images by computing the FID, CLIP-FID, and IS of the AI-generated image set and the referenced natural image set. The results of these set-level measures can reflect the difference between AIGIs and natural ones from an integral perspective.
Based on the different levels of image quality measures, our framework can assess and measure the fine-grained AI-Natural images differences with Full-reference and Non-reference angles, covering both low-level and high-level. More details about the utilized traditional image quality assessment measures are described in Appendix C.
@@ 2. Semantic Alignment Discrepancy:
Generative models require the referenced prompts as semantic guidance to generate images. Therefore, semantic alignment can be an important quality indicator of the AIGIs. We use the widely used CLIP model to calculate the difference rates of CLIP Scores (Hessel et al., 2021), $DR_{CS}$ to represent alignment discrepancies between AIGIs and natural images, and provide a fine-grained analysis with respect to both unimodal and multimodal guidance.